OpenStudy (anonymous):

Solve this system using the substitution method. 3x + 2y = 10 2x + 3y = 15/2 Answer No solution y = (-3/2)x + 5 x = 3, y = -1/2 x = 3, y = 1/2

OpenStudy (zpupster):

get y from 1st equation and substitute in 2nd eq 3x + 2y = 10 subtract 3x both sides 2y= 10-3x now divide both sides by 2 2y/2=(10-3x)/2 y=10/2 -3x/2 or y= 5 -3/2x now we know y so sub it to 2nd eq 2x + 3(5 -3/2x) = 15/2 mult thru by 3 to get rid of parentheses 2x + 15 -9/2x=15/2 lets get a common denominator of 2 4/2x + 30/2 -9/2x = 15/2 subtract 30/2 both sides 4/2x-9/2x=-15/2 now combine x terms -5/2x = -15/2 mult both sides by -2/5 x=30/10 or = 3 x = 3 now plug it in both equations and solve for y also for this step 2x + 15 -9/2x=15/2 i could just mult by 2 and that would get rid of the denominator 4x+30-9x = 15 solve you get the same